OpenStudy (anonymous):

Given the two polygons are similar, calculate the value of x. 4.1 3.98 3.89 2.98

OpenStudy (mrnood):

when figures are 'similar' it means they have the same 'proprtions' so all angles ar eth esame, and all ratios of edges are the same So look at the big rectancle and write the 'ratio' of th small edge to the large edge Then do the same for the small rectangle (this time you need to leave x in the proportion' these two ratios are equal to each otherso you can work out x from that

OpenStudy (anonymous):

I think its 2.98 @MrNood

OpenStudy (mrnood):

do oyu have a reason for thinking that? show your working - or work through my suggestion (hint - I think you need to...)

OpenStudy (mrnood):

the ratio od the small edge to the large edge is \[\frac{ smalledge }{ longedge}\] write that ration for both rectangels

OpenStudy (anonymous):

cause you subtracted 4 from 9 to get 5 and then do the same to the 7 and 2.98 is the closest to 3

OpenStudy (mrnood):

this is not about subtraction - it is about scale factors - which means division or multiplication try doing th ethings I suggested above...

OpenStudy (anonymous):

5/9=0.56

OpenStudy (anonymous):

and x/7 @MrNood

OpenStudy (mrnood):

yes - thoae are correct for the 2 ratios now because the rectangles are similar those ratios are equal to each other so write them as en equation (ratio 1 = ratio 2) and then find what x is..

OpenStudy (anonymous):

5/9=x/7 i forget how to cross multiply @MrNood

OpenStudy (anonymous):

nevermind i got 3.89